OpenStudy (anonymous):

choose a point-slope form of the question below that represents the line that passes through the point (-1,6) and has a slope of -3

OpenStudy (kidrah69):

We need to start off with the point slope formula y-y1=m(x-x1)

OpenStudy (kidrah69):

Just plug these in :) y1=6 slope=m=-3 x1=-1

OpenStudy (anonymous):

is there any answer choices

OpenStudy (anonymous):

here are the answer choices i forgot to put. A. y-6 =-3X-3 B. y-6 = -3(x+1) C. y = -3x +3 D. 3x +y = 3

OpenStudy (kidrah69):

Right i gave you the formula and the values to plug in now just plug them in :)

OpenStudy (anonymous):

i'm gonna say...........B

OpenStudy (kidrah69):

yep :)

OpenStudy (kidrah69):

Good job :) also \(\huge\color{red}{Welcome~to~Openstudy}\)

OpenStudy (anonymous):

yay!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

thank you so much!!!!!!!

OpenStudy (kidrah69):

:)